OpenStudy (anonymous):

Determine whether each function is even, odd, or neither: y= x^3 sinx

OpenStudy (unklerhaukus):

\[y=x^3 \sin x\] y is the product of an odd function and an odd function is even function

OpenStudy (anonymous):

thus function is even because y=(-x)^3*sin(-x) = -x^3(-sinx) = x^3(sinx) therefore f(x) = f(-x)